OpenStudy (anonymous):

What is the distance between P(0, 14) and Q(5, -2)?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf distance~=~\sqrt{(y_1-y_2)^{2}+(x_1 - x_2)^{2}} }\)

OpenStudy (solomonzelman):

plug in your coordinates into the above formula, as x1=0 x2=5 y1=14 y2=-2

OpenStudy (anonymous):

I putt that formula in my TI-nspire and it gave me 7.2111 but my options are A the square root of 281 B 13 C the square root of 171 D 14

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt{(y_1-y_2)^{2}+(x_1 - x_2)^{2}} }\) \(\Large\color{blue}{ \bf \sqrt{(5-(-2)~~)^{2}+(0 -14)^{2}} }\) \(\Large\color{blue}{ \bf \sqrt{(5+2)^{2}+(0 -14)^{2}} }\) \(\Large\color{blue}{ \bf \sqrt{(7)^{2}+( -14)^{2}} }\) \(\Large\color{blue}{ \bf \sqrt{49+196} }\) \(\Large\color{blue}{ \bf \sqrt{245} }\)